OpenStudy (anonymous):

Simplify the following rational function: 1/x-1/x-1

OpenStudy (chaise):

Depends what you call simple: 1/(x-2)^2 or (x-2)^-2

OpenStudy (chaise):

(x-1)'s instead of (x-2) sorry D:

OpenStudy (anonymous):

Assuming you mean for the denomintor of the second fraction to be x-1 you would have to get a LCD of x(x-1)\[1/x-1/(x-1)=(x-1-x)/(x(x-1)=-1/(x^2-x)\]

OpenStudy (anonymous):